OpenStudy (anonymous):

make a conjecture about the next item in the sequence 6,8,-32,-30,120

OpenStudy (anonymous):

6 + 2 = 8 8 * (-4) = -32 -32 + 2 = -30 -30 * (-4) = 120 Do you see the pattern?

OpenStudy (anonymous):

Kindaa

OpenStudy (anonymous):

So, what do you tink the next number is?